OpenStudy (anonymous):

Which of the following represents this function written in intercept form? y= -x^2 - x +6 a. y= -(x+2)(x+6) b. y= -(x-2)(x+3) c. y= (-x+3)(x-4) d. y= -(x+2)(x-6)

OpenStudy (anonymous):

is b the right answer?

OpenStudy (anonymous):

Are you trying to factor it? How do you write a quadratic equation in the form of intercept form?

OpenStudy (anonymous):

y=a(x-r1)(x-r2)

OpenStudy (aum):

-x^2 - x + 6 = -{ x^2 + x - 6 } = -{ x^2 + 3x - 2x - 6 } = -{ x(x+3) - 2(x+3) } = -(x-2)(x+3)